# Relative-error monotonicity testing

### v1

@@ 6.1 Distributions $\mathcal{D}_{\mathrm{yes}}$ and $\mathcal{D}_{\mathrm{no}}$
    $\bm{\Gamma}_{\bm{T}}(x)=\begin{cases}0^{*}&T_{i}(x)=0\ \text{for all $i\in[L]$}\\ 1^{*}&T_{i}(x)=1\ \text{for at least two different $i\in[L]$}\\ i&T_{i}(x)=1\ \text{for a unique $i\in[L]$}.\end{cases}$
@@
    Sample ${\bm{H}}=({\bm{h}}_{i}\colon i\in[L])$ from $\mathcal{E}_{\text{yes}}$, where each ${\bm{h}}_{i}\colon\{0,1\}^{n}\to\{0,1\}$ is independently (1) with probability $2/3$, a random dictatorship Boolean function, i.e., ${\bm{h}}_{i}(x)=x_{k}$ with $k$ sampled uniformly at random from $[n]$; and (2) with probability $1/3$, the all-00 function.
@@
    ${\bm{f}}(x)=\begin{cases}0&\bm{\Gamma}_{{\bm{T}}}(x)=0^{*}\\ 1&\bm{\Gamma}_{{\bm{T}}}(x)=1^{*}\\ {\bm{h}}_{\bm{\Gamma}(x)}(x)&\text{otherwise (i.e., $\bm{\Gamma}_{{\bm{T}}}(x)\in[L]$)}\end{cases}$
@@
This finishes the proof of 22. ∎
Note that the number of points on the $43n/$ 4 layer is (${n\choose 3n/4}\geq|f^{-1}(1)|/100$. Thus, the expected size of $\bm{X}$ is at least $|f^{-1}(1)|/10000$. Thus, we have $|\bm{X}|=\Omega(|f^{-1}(1)|)$ with probability at least 0.0001. This finishes the proof of Lemma 21. ∎
Let $\varepsilon_{0}$ and $c$ be the two constants hidden in the $\Omega(1)$’s in Lemma 21, i.e., a function ${\bm{f}}\sim\mathcal{D}_{\mathrm{no}}$ has relative distance to monotonicity at least $\varepsilon_{0}$ with probability at least $c$. Let $\alpha$ be the constant $\min(\varepsilon_{0},c)/100$. Let $q=n/(C_{0}\log n)$ for some sufficiently large constant $C_{0}$ to be specified later. We show that no randomized, $q$-query, non-adaptive, $\mathrm{MQ}$-only algorithm can distinguish monotone functions from functions that have relative distance at least $\varepsilon_{0}$ with probability at least $1-\alpha$. This implies (using standard amplification techniques) that any randomized, non-adaptive, $\mathrm{MQ}$-only algorithm that distinguishes the two cases with probability $(99/100)\cdot 2/3$ must make $\Omega(q)$ queries as desired.
To this end, we use Yao’s minimax principle and consider a deterministic, $q$-query, non-adaptive, $\mathrm{MQ}$-only algorithm ALG which consists of a set $Q$ of $q$ query strings $x_{1},\dots,x_{q}$ and a decision procedure which outputs “accept” or “reject” given $f(x_{k})$ for each $k\in[q]$. In order to prove the lower bound, it suffices to prove the following lemma:
@@
Let ALG be any non-adaptive, deterministic, $q$-query algorithm. Then
$\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{yes}}}\big{[}\text{ALG}\text{ accepts }{\bm{f}}\big{]}\leq(1+\alpha)\cdot\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{no}}}\big{[}\text{ALG}\text{ accepts }{\bm{f}}\big{]}+\alpha.$
Assume Lemma 23, and assume for a contradiction that there is a randomized, nonadaptive, $q$-query algorithm that can distinguish monotone functions from functions that have relative distance at least $\varepsilon_{0}$ to $\mathcal{C}_{\textsf{monotone}}$ with probability at least $1-\alpha$. This implies a non-adaptive, deterministic algorithm ALG such that
$\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{yes}}}\big{[}\text{ALG}\text{ accepts }{\bm{f}}\big{]}-\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{no}}}\big{[}\text{ALG}\text{ accepts }{\bm{f}}\big{]}\geq(1-\alpha)-(1-c(1-\alpha))\geq c/2,$
using the choice of $\alpha$. This contradicts Lemma 23, again using the choice of $\alpha$.
@@ 6.2 Proof of Lemma 23
Given that Lemma 23 is only about running ALG on $\mathcal{D}_{\mathrm{yes}}$ and $\mathcal{D}_{\mathrm{no}}$, we may assume without loss of generality that every point queried by ALG lies in the two layers $43n/$ 4 and $13n/4+$ 1. Given a function $f=f_{T,H}$ from the support of either $\mathcal{D}_{\mathrm{yes}}$ or $\mathcal{D}_{\mathrm{no}}$ and a set $Q$ of $q$ queries, we define the following tuple $(P;R;\rho)$ as the outcome of $Q$ on $f$, where $P=(P_{i}\mathrel{\mathop{\ordinarycolon}}i\in[L])$ with $P_{i}\subseteq Q$, $R=(R_{i}\mathrel{\mathop{\ordinarycolon}}i\in[L])$ with $R_{i}\subseteq Q$, and $\rho=(\rho_{i}\mathrel{\mathop{\ordinarycolon}}i\in[L])$ with $\rho_{i}\mathrel{\mathop{\ordinarycolon}}P_{i}\rightarrow\{0,1\}$:
@@
    We build $P$ and $R$ as follows. For each $x\in Q$ (the order does not matter), if no term in $T$ is satisfied, add $x\rightarrow R_{i}$ for all $i\in[L]$; if $x$ satisfies a unique term in $T$, say $T_{i}$, add $x\rightarrow P_{i}$ and $x\rightarrow R_{j}$ for all $j\neq i$; if $x$ satisfies more than one term in $T$, letting $i<i^{\prime}$ be the first two terms satisfied by $x$, add $x\rightarrow P_{i}$, $x\rightarrow P_{i^{\prime}}$, and $x\rightarrow R_{j}$ for all $j\mathrel{\mathop{\ordinarycolon}}j<i^{\prime}$ and $j\neq i$.
@@
    For each $i\in[L]$ such that $P_{i}\neq\emptyset$ and $x\in P_{i}$, set $\rho_{i}(x)=h_{i}(x)$.
We record the following facts about outcomes of $Q$ on functions in the support of $\mathcal{D}_{\mathrm{yes}}$ and $\mathcal{D}_{\mathrm{no}}$, which are evident by construction:
@@
The $P$ and $R$ parts of the outcome depend on $T$ only. Once $P$ is fixed, the $\rho$ part of the outcome depends on $H$ only.
Moreover, values of $f(x)$, $x\in Q$, can be determined by the outcome of $Q$ on $f$. As a result, if the outcomes of $Q$ on two functions $f$ and $f^{\prime}$ are the same, then ALG returns the same answer.
Fact 25.
$\sum_{i\in[L]}|P_{i}|\leq 2q$.
The following definition of bad vs good outcomes is crucial to the analysis:
@@
Let $(P;R;\rho)$ be a possible outcome of $Q$ on a function $f$ from the support of $\mathcal{D}_{\mathrm{yes}}$ and $\mathcal{D}_{\mathrm{no}}$. We say $(P;R;\rho)$ is a bad outcome if at least one of the following two events happen:
@@
    |$\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=y_{k}=1\}\big{|}\leq 3n/4-100\log n.$
@@
Let $Q$ be a set of $q$ queries. Then we have
$\mathop{\operatorname{{\bf Pr}}}_{{\bm{T}}\sim\mathcal{E},{\bm{H}}\sim\mathcal{E}_{\text{yes}}}\big{[}\hskip 0.85355pt\text{outcome of $Q$ on $f_{{\bm{T}},{\bm{H}}}$ is bad}\hskip 0.85355pt\big{]}\leq\alpha.$
@@
Let $Q$ be a set of $q$ queries. For any good outcome $\ell=(P;R;\rho)$, we have
$\mathop{\operatorname{{\bf Pr}}}_{{\bm{T}}\sim\mathcal{E},{\bm{H}}\sim\mathcal{E}_{\text{yes}}}\big{[}\hskip 0.85355pt\text{outcome of $Q$ on $f_{{\bm{T}},{\bm{H}}}$ is $\ell$}\hskip 0.85355pt\big{]}\leq(1+\alpha)\cdot\mathop{\operatorname{{\bf Pr}}}_{{\bm{T}}\sim\mathcal{E},{\bm{H}}\sim\mathcal{E}_{\text{no}}}\big{[}\hskip 0.85355pt\text{outcome of $Q$ on $f_{{\bm{T}},{\bm{H}}}$ is $\ell$}\hskip 0.85355pt\big{]}.$
Proof of Lemma 23 assuming 27 and 28.
By 24 the outcome of $Q$ on $f$ uniquely determines $f(x)$, $x\in Q$. This allows us to define $L_{1}$ as the set of outcomes such that ALG accepts $f$ if and only if the outcome of $Q$ on $f$ lies in $L_{1}$.
By 27 and 28, we have
$\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}}\big{[}\text{ALG}\ \text{accepts}\ {\bm{f}}\big{]}$ $=\sum_{\ell\in L_{1}}\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}}\big{[}\text{outcome of $Q$ on ${\bm{f}}$ lies in $L_{1}$}\big{]}$
$\leq\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}}\big{[}\text{outcome of $Q$ on ${\bm{f}}$ is bad}\big{]}$
$\hskip 28.45274pt+\sum_{\text{good $\ell\in L_{1}$}}\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}}\big{[}\text{outcome of $Q$ on ${\bm{f}}$ is $\ell$}\big{]}$
$\leq\alpha+\sum_{\text{good $\ell\in L_{1}$}}(1+\alpha)\cdot\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{no}}}\big{[}\text{outcome of $Q$ on ${\bm{f}}$ is $\ell$}\big{]}$
$\leq(1+\alpha)\cdot\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{no}}}\big{[}\text{ALG}\ \text{accepts}\ {\bm{f}}\big{]}+\alpha.$

### v2

@@ 5.2 Distributions $\mathcal{D}_{\mathrm{yes}}$ and $\mathcal{D}_{\mathrm{no}}$
[…]
@@
    Sample ${\bm{H}}=({\bm{h}}_{i}\colon i\in[L])$ from $\mathcal{E}_{\text{yes}}$, where each ${\bm{h}}_{i}\colon\{0,1\}^{n}\to\{0,1\}$ is independently (1) with probability $2/3$, a random dictatorship Boolean function, i.e., ${\bm{h}}_{i}(x)=x_{k}$ with $k$ sampled uniformly at random from $[n]$; and (2) with probability $1/3$, the all-0 function.
@@
    ${\bm{f}}(x)=\begin{cases}0&\bm{\Gamma}_{{\bm{T}}}(x)=0^{*}\\ 1&\bm{\Gamma}_{{\bm{T}}}(x)=1^{*}\\ {\bm{h}}_{\bm{\Gamma}(x)}(x)&\text{otherwise (i.e., $\bm{\Gamma}_{{\bm{T}}}(x)\in[L]$).}\end{cases}$
@@
This finishes the proof of 19. ∎
Note that the number of points on the $3n/4$ layer is
${n\choose 3n/4}\geq\frac{|f^{-1}(1)|}{100}.$
Thus, the expected size of $\bm{X}$ is at least $|f^{-1}(1)|/10000$. Thus, we have $|\bm{X}|=\Omega(|f^{-1}(1)|)$ with probability at least 0.0001. This finishes the proof of Lemma 18. ∎
Let $\varepsilon_{0}$ and $c$ be the two constants hidden in the $\Omega(1)$’s in Lemma 18, i.e., a function ${\bm{f}}\sim\mathcal{D}_{\mathrm{no}}$ has relative distance to monotonicity at least $\varepsilon_{0}$ with probability at least $c$. Let $\alpha=\min(\varepsilon_{0},c)/100$.
We will consider deterministic, $q$-query, (two-stage) non-adaptive, $\mathrm{MQ}$-only algorithms ALG, which are defined as follows.
Definition 20.
A deterministic, $q$-query, (two-stage) non-adaptive, $\mathrm{MQ}$-only algorithms ALG runs on $\mathcal{D}_{\mathrm{yes}}$ and $\mathcal{D}_{\mathrm{no}}$ as follows. Upon an input function $f$ from either $\mathcal{D}_{\mathrm{yes}}$ or $\mathcal{D}_{\mathrm{no}}$,
  1. 1.
    In the first stage, ALG draws a sequence of $q$ samples $\bm{S}$ independently and uniformly at random from the set of all points at or above layer $3n/4$, and queries all of them.
  2. 2.
    In the second stage, based on $\bm{S}$ and the query results (denoted by $f(\bm{S})$, a $q$-bit string), ALG picks deterministically a set $\mathbf{Q}=Q(\bm{S},f(\bm{S}))$ of $q$ points to query, and finally accepts or rejects based on all the information in $(\mathbf{Q},\bm{S},f(\mathbf{Q}\cup\bm{S}))$.
In order to prove Theorem 14, it suffices to prove the following lemma:
@@
Let $q=n/(C_{0}\log n)$ for some sufficiently large constant $C_{0}$. For any deterministic, $q$-query, (two-stage) non-adaptive, $\mathrm{MQ}$-only algorithm ALG, we have
$\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{yes}}}\big{[}\text{ALG}\text{ accepts }{\bm{f}}\big{]}\leq(1+\alpha)\cdot\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{no}}}\big{[}\text{ALG}\text{ accepts }{\bm{f}}\big{]}+2\alpha.$
To see that this proves the lower bound in Theorem 14, assume for a contradiction that there is a randomized, $o(q)$-query, non-adaptive algorithm that accepts every monotone function with probability at least $1-\alpha$ and rejects every function that has relative distance at least $\varepsilon_{0}$ to monotonicity with probability at least $1-\alpha$. Then by Yao’s minimax principle, there exists a deterministic, $q$-query, (standard) non-adaptive algorithm $\text{ALG}^{\prime}$ that satisfies
$\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{yes}}}\big{[}\text{ALG}^{\prime}\text{ accepts }{\bm{f}}\big{]}-\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{no}}}\big{[}\text{ALG}^{\prime}\text{ accepts }{\bm{f}}\big{]}\geq(1-\alpha)-\big{(}1-c(1-\alpha)\big{)}\geq c\big{/}2.$
Note that given any input function $f$ from either $\mathcal{D}_{\mathrm{yes}}$ or $\mathcal{D}_{\mathrm{no}}$, such an algorithm $\text{ALG}^{\prime}$ works by first drawing a sequence of $o(q)$ samples from $f^{-1}(1)$ and then, based on the samples drawn, picking deterministically a set of $o(q)$ points to query. Using 15, with a suitably small choice of the constant $\tau>0$, we can obtain from $\text{ALG}^{\prime}$ a deterministic, $o(q)$-query, (two-stage) non-adaptive, $\mathrm{MQ}$-only algorithm ALG that satisfies
$\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{yes}}}\big{[}\text{ALG}\text{ accepts }{\bm{f}}\big{]}-\mathop{\operatorname{{\bf Pr}}}_{{\bm{f}}\sim\mathcal{D}_{\text{no}}}\big{[}\text{ALG}\text{ accepts }{\bm{f}}\big{]}\geq c/4,$
which, however, contradicts with Lemma 21 using the choice of $\alpha$.
@@ 5.3 Proof of Lemma 21
Let ALG be a deterministic, $q$-query, (two-stage) non-adaptive, MQ-only algorithm in the rest of the section, where the parameter $q=n/(C_{0}\log n)$ for some sufficiently large constant $C_{0}$. We start with a definition that gives the outcome of an $f$ from either $\mathcal{D}_{\mathrm{yes}}$ or $\mathcal{D}_{\mathrm{no}}$ on a set $V$ of query points.
Definition 22.
Let $f=f_{T,H}$ be a function from the support of either $\mathcal{D}_{\mathrm{yes}}$ or $\mathcal{D}_{\mathrm{no}}$, and $V$ be a set of points that lie at or above layer $3n/4$. We define the following tuple $(P;R;\rho)$ as the outcome of $f$ on $V$, where $P=(P_{i}\mathrel{\mathop{\ordinarycolon}}i\in[L])$ with $P_{i}\subseteq V$, $R=(R_{i}\mathrel{\mathop{\ordinarycolon}}i\in[L])$ with $R_{i}\subseteq V$, and $\rho=(\rho_{i}\mathrel{\mathop{\ordinarycolon}}i\in[L])$ with $\rho_{i}\mathrel{\mathop{\ordinarycolon}}P_{i}\rightarrow\{0,1\}$ for each $i\in[L]$:
@@
    We build $P$ and $R$ as follows. For each $x\in V$ that lies in layer $3n/4$ or $3n/4+1$ (the order does not matter), if no term in $T$ is satisfied, add $x\rightarrow R_{i}$ for all $i\in[L]$; if $x$ satisfies a unique term in $T$, say $T_{i}$, add $x\rightarrow P_{i}$ and $x\rightarrow R_{j}$ for all $j\neq i$; if $x$ satisfies more than one term in $T$, letting $i<i^{\prime}$ be the first two terms satisfied by $x$, add $x\rightarrow P_{i}$, $x\rightarrow P_{i^{\prime}}$, and $x\rightarrow R_{j}$ for all $j\mathrel{\mathop{\ordinarycolon}}j<i^{\prime}$ and $j\neq i$.
@@
    Finally, for each $i\in[L]$ with $P_{i}\neq\emptyset$, set $\rho_{i}(x)=h_{i}(x)$ for each $x\in P_{i}$.
Observe that having $x\in P_{i}$ indicates that $x$ satisfies term $T_{i}$ (though not every $x$ that satisfies a term $T_{i}$ will be in $P_{i}$), and that having $x\in R_{i}$ indicates that $x$ does not satisfy term $T_{i}$ (though not every $x$ that does not satisfy term $T_{i}$ will be in $R_{i}$).
We will denote the outcome $(P;R;\rho)$ of $f$ on $V$ as $\mathsf{outcome}(f,V)$.
We record the following facts that are evident by construction:
Fact 23.
The $P$ and $R$ parts of the outcome depend on $T$ only. Once $T$ is fixed, the $\rho$ part of the outcome depends on $H$ only. Moreover, values of $f(x)$, $x\in V$, can be determined from $\mathsf{outcome}(f,V)$.
@@
$\sum_{i\in[L]}|P_{i}|\leq 2|V|$.
Let $\mathcal{H}_{\text{yes}}$ denote the following distribution: To draw a tuple $\bm{U}=(\mathbf{Q},\bm{S},(\mathbf{P};\bm{R};\bm{\rho}))\sim\mathcal{H}_{\text{yes}}$, we first draw ${\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}$ and a sequence of $q$ points $\bm{S}$ uniformly at random from points that are at or above layer $3n/4$. $\mathbf{Q}$ is then the set of $q$ points that ALG queries in the second stage, given $\bm{S}$ and ${\bm{f}}(\bm{S})$ in the first stage, and $(\mathbf{P};\bm{R};\bm{\rho})$ is set to be $\mathsf{outcome}({\bm{f}},\mathbf{Q}\cup\bm{S})$. We define $\mathcal{H}_{\text{no}}$ similarly, with the function ${\bm{f}}\sim\mathcal{D}_{\mathrm{no}}$ instead of $\mathcal{D}_{\mathrm{yes}}$.
@@
Let $U=(Q,S,(P;R;\rho))$ be a tuple in the support of either $\mathcal{H}_{\text{yes}}$ or $\mathcal{H}_{\text{no}}$. We say $U$ is bad if at least one of the following two events happen:
@@
    $\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=y_{k}=1\}\big{|}\leq 3n/4-100\log n.$
@@
A tuple $\bm{U}\sim\mathcal{H}_{\text{yes}}$ is bad with probability at most $2\alpha$.
@@
For any fixed good tuple $U$ in the support of $\mathcal{H}_{\text{yes}}$, we have
$\mathop{{\bf Pr}\/}_{\bm{U}\sim\mathcal{H}_{\text{yes}}}\big{[}\hskip 0.85355pt\bm{U}=U\hskip 0.85355pt\big{]}\leq(1+\alpha)\cdot\mathop{{\bf Pr}\/}_{\bm{U}\sim\mathcal{H}_{\text{no}}}\big{[}\hskip 0.85355pt\bm{U}=U\hskip 0.85355pt\big{]}.$
Before proving 26 and 27, we use them to prove Lemma 21:
Proof of Lemma 21 assuming 26 and 27.
Let $\mathcal{U}$ be the set of those $U$ in the support of $\mathcal{H}_{\text{yes}}$ that lead ALG to accept. Using 26 and 27, we have
$\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}}\big{[}\text{ALG}\ \text{accepts}\ {\bm{f}}\big{]}$ $=\sum_{U\in\mathcal{U}}\mathop{{\bf Pr}\/}_{\bm{U}\sim\mathcal{H}_{\text{yes}}}\big{[}\bm{U}=U\big{]}$
$\leq\sum_{\text{good}\ U\in\mathcal{U}}\mathop{{\bf Pr}\/}_{\bm{U}\sim\mathcal{H}_{\text{yes}}}\big{[}\bm{U}=U\big{]}+2\alpha$ (4)
$\leq(1+\alpha)\sum_{\text{good}\ U\in\mathcal{U}}\mathop{{\bf Pr}\/}_{\bm{U}\sim\mathcal{H}_{\text{no}}}\big{[}\bm{U}=U\big{]}+2\alpha$ (5)
$\leq(1+\alpha)\cdot\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{no}}}\big{[}\text{$\text{ALG}$ accepts ${\bm{f}}$}\big{]}+2\alpha,$
where Equation 4 used 26 and Equation 5 used 27, and the last inequality is because
$\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{no}}}\big{[}\text{$\text{ALG}$ accepts ${\bm{f}}$}\big{]}=\sum_{\text{good}\ U\in\mathcal{U}}\mathop{{\bf Pr}\/}_{\bm{U}\sim\mathcal{H}_{\text{no}}}\big{[}\bm{U}=U\big{]}+\sum_{\text{bad}\ U\in\mathcal{U}}\mathop{{\bf Pr}\/}_{\bm{U}\sim\mathcal{H}_{\text{no}}}\big{[}\bm{U}=U\big{]}.$